What is the MSE for a simple linear regression with SST = 150, SSR = 100, and n = 12? 5 15 10 4.5
Answer #1

SST is the total sum of squares

SSE is the sum of the square of errors

SSR = SST -SSE gives SSE = SST - SSR = 150-100 = 50

Mean square error = SSE/(n-2) = 50/(12-2) = 50/10 = 5

Hence correct answer is 5.